OpenStudy (anonymous):

2 C2H6 + 7 O2 yields 4 CO2 + 6 H2O What volume of ethane is needed to react with 21L of oxygen? T and P have already been calculated

OpenStudy (somy):

what are the values for P and T?

OpenStudy (anonymous):

it doesn't say. it just says they have been calculated so I am assuming they must be constant?

OpenStudy (somy):

yeah that means they just want to confuse you

OpenStudy (somy):

is there anymore info? like if the reaction was carried out under room temp and pressure or standard temp and pressure

OpenStudy (somy):

i think most probably its STP

OpenStudy (somy):

what do u think?

OpenStudy (anonymous):

ye i think it must be STP, but this was found in the law of combining volumes section.

OpenStudy (somy):

anyways the formula u use here is \[volume = no. of moles \times STP\]

OpenStudy (somy):

are u familiar with this formula?

OpenStudy (anonymous):

not really. so should i convert the 21:L O2 to ethane mole and multiply by the STP?

OpenStudy (somy):

nah first u find mole of 21 L of O2

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

so convert L to grams then find moles of O2 then convert to ethane then multiply by stp?

OpenStudy (somy):

no of mole= volume/STP = 21/22.4= 0.9375 mole

OpenStudy (somy):

nooo forget about grams u fdon't even need grams

OpenStudy (somy):

don't*

OpenStudy (anonymous):

hmmmm ok, but the answer is 6L

OpenStudy (somy):

yeah wait lol

OpenStudy (somy):

so now u know the mole of O2 and now look at the ratio between O2 and C2H6

OpenStudy (somy):

its 2:7 right?

OpenStudy (somy):

so if 7---- 0.9375 moles 2---- ? so do cross multiplication and find the mole of C2H6

OpenStudy (somy):

are u getting it?

OpenStudy (somy):

anyways do this you'll get the mole since you do it in calculator, do this 0.9375*2/7= .....mole and do not round off and straightaway multiply it by 22.4 = and u'll get 6 L

OpenStudy (somy):

btw here is the link for proving that its STP http://chemistry.tutorvista.com/inorganic-chemistry/law-of-combining-volumes.html

OpenStudy (anonymous):

ok thanks

OpenStudy (somy):

did u get it though?

OpenStudy (anonymous):

yes. what I did not know was the #moles=volume/STP

OpenStudy (somy):

yeah good then :) happy to help :D